OpenStudy (anonymous):

3ln(x^2)-2ln(x 1/3) +4ln(x-1)=1 ,solve for x: anyone?

OpenStudy (anonymous):

Is the problem\[3\ln(x^{2})-2\ln(x^{\frac{1}{3}})+4\ln(x-1)=1\]?

OpenStudy (anonymous):

yeah@alex

OpenStudy (anonymous):

do you need real solutions or all solutions? exact solutions?

OpenStudy (anonymous):

Real Solution in Exact Form: x = 1/6 (3+sqrt(3 (3+4 sqrt(3) e^(1/4))))

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

wait is that x to the 1/3

OpenStudy (anonymous):

\[3lnx^2-2lnx^{\frac{1}{3}}+4\ln(x-1)=1\]\[lnx^{6}-lnx^{\frac{2}{3}}+\ln(x-1)^4=1\]\[\ln \frac{x^6}{x^{\frac{2}{3}}}+ln(x-1)^4=1\]\[\ln x^{\frac{16}{3}}+\ln(x-1)^4=1\]\[\ln[x^{\frac{16}{3}}\times(x-1)^4]=lne\]\[[x^{\frac{16}{3}}\times(x-1)^4]=e\]I'm too lazy to complete it, sorry.

OpenStudy (kira_yamato):

I'll leave you with the basic logarithm from here...

OpenStudy (anonymous):

thnkuu so much kira..

OpenStudy (anonymous):

Kira - How did you make a picture like that?

OpenStudy (kira_yamato):

Screenshot...

OpenStudy (anonymous):

Of what? MathType, or similar? Sorry for being inquisitive, but that was nice. xD

OpenStudy (kira_yamato):

I used equation editor... But you can use MathType if you like... It's the same...